OpenStudy (yuno._.gasai):

what is the solution to y + 7 = -9

OpenStudy (yuno._.gasai):

what is the solution to 8 * g = -32

OpenStudy (kweenlexis):

Hold up right quick

OpenStudy (yuno._.gasai):

what is the solution to -11 * k= -33

OpenStudy (yuno._.gasai):

kay

OpenStudy (kweenlexis):

It's -4 for 8*g= -32 because 8* (-4)=-32

OpenStudy (kweenlexis):

Do you have any more?

OpenStudy (yuno._.gasai):

the other two

OpenStudy (kweenlexis):

what are them?

OpenStudy (yuno._.gasai):

what is the solution to -11 * k= -33 what is the solution to y + 7 = -9

OpenStudy (kweenlexis):

-11 * k =-33 k=3 y+7=-9 y=-16

OpenStudy (yuno._.gasai):

thank you

OpenStudy (kweenlexis):

No problem, my bad im so late.